OpenStudy (anonymous):

Find the coordinates for the center and length of radius for this circle. (x - 4)2 + y2 = 12 Center = (_,_) Radius = radius _sqrt_

OpenStudy (waleed_imtiaz):

Lets first expand the equation to convert it into the general form of the equation i.e. x^2+y^2+2gx+2fy+c=0 So on expansion it will become, x^2+16-8x+y^2=12 now, x^2+y^2-8x+4=0 now centre will be (-g,-f) ... compare the expanded equation with the general equation I wrote above... U will get centre (4,0) ... c=4 and g=-4, f=0 now the formula for the length of the radius is= \[\sqrt{g^{2}+f^{2}-c} \] put the values,,, U will get \[\sqrt{12}\]

OpenStudy (waleed_imtiaz):

Got it @Sephichief , any problem, can ask me ........ thank you